Home‎ > ‎Crew 147‎ > ‎

115-CommandersReport

Commander report
01/15/2015
Romain Charles

We started our day with the promise of a new EVA under the sun. It's
always good for the moral.

EVA - Wissam (EVA Lead), Rebeca, Scott, Romain
Like yesterday, the debriefings that we had together helped to work
efficiently during the EVA preparation. We could tests a draft "EVA
Checklist" which will  help each EVA Lead to be ready and efficient.
The EVA lasted from 10:10 to 12:40 and here are a few highlights:
• Engineering activity to review the status of our external systems
• Quadcopter experiment in 2 different locations
Upon return to the Hab, we did a debriefing of this 3rd EVA and one of the
main topic was the high rate of communication issues.

Communication issues
Soon after our lunch we started an extensive check of all our
communication devices (radios & headsets) in all the different
configurations. You will receive more details about the testing in
Rebeca's Engineering report but the conclusion is that the microphones of
3 headsets (out of 6) are not working.
From our different tests, we now assume that the humidity of the helmet
could be the cause of this high rate of failure amongst the microphones.
We would like to request a near-future resupply of EVA microphone
equipment with a minimum of 3 new headsets so we can continue our
exploration of the Martian surface safely.

Engineering EVA
The ATV situation was evaluated again during the Engineering activities of
our morning EVA but to no avail. An Engineering EVA was then planned from
17h30 to 18h to have a clear understanding of the situation. Rebeca was
able to start 4 ATVs (out of 5) with success but the ATV5 is still a
problem. You will find all the details in her Engineering report.

Water contingency
We monitored the Static tank status during our EVA's Engineering
activities at 10:30 and also during the Engineering EVA at 18:10 but it
didn't change. The tank is still low with water. Our crew continues to
follow the "Water saving" scenario but, as explained on the 13th of
January, tomorrow is the last day when we have enough water supplies.

We urgently request the status of the water resupply. Shall we schedule a
"water recovery" activity during our planned EVA tomorrow?

Upcoming EVA
We are sending a request for approval of an EVA on Jan 16th 2015. As the
weather conditions improved for the last couple of days, Danielle (EVA
Lead), Andrew, Renee and Scott are planning an EVA to conduct a test of
the Phoenix Rover.

I finish this report a bit abruptly as, for the 3rd day in a row, we
caught a glimpse of our new mascot, Marvin the mouse, and we're starting a
hunt!